Harvard-Smithsonian Center for Astrophysics
FLWO Ridge Telescopes
Issue No. 6, February 2006
The present memo is another report from the F. L. Whipple
Observatory. We chronicle FLWO activities and the status and progress
of telescopes, instruments and related equipment. The telescopes
involved are the 1.5m Tillinghast, the 1.2m, the 1.3m PAIRITEL and the
HATs. This belated report covers the period Apr 2005-Feb 2006. All
issues of these reports are posted at
If you have any comments, or for further information on any of the
projects below, please contact us.
The instruments currently available are the Keplercam CCD (dark
time) and the stelircam IR camera (bright time).
- The FTS chiller system for the primary mirror has been in operation
since Spring 2003. The system maintains the mirror temperatures
within about 0.2C of the telescope chamber ambient temperature under
observing conditions (see the link to temperatures on the 1.2m web
page). During Spring 2005, we found contamination in the glycol lines
and a small leak, accompanied by slower temperature slews than in
previous seasons. RH replaced the hoses, fittings and glycol. The new
hoses are expected to be impervious to degradation by glycol. The
system returned to its initial high efficiency in late September
2005. Unfortunately, in Oct 2005 RH found a new significant glycol
leak in the chiller where none had been seen before. That required
draining the system. RH also found that the glycol pump needed to be
rebuilt. The chiller went back in operation in Dec 2005. However, its
efficiency was nowhere near its achievable level. RH discovered that
the heat exchangers attached to the mirror cell were clogged in an
irreversible fashion. RH acquired replacement heat exchangers and
replaced the glycol coolant with a flavor friendlier to the Al in the
exchangers. The chiller has been operating at high efficiency since
early Feb 2006.
- Remote observing remains popular. Observers who plan to use it
should ensure to read the documentation and to contact the staff well
before their scheduled runs. It is essential for observers who have
never conducted remote observing, who are changing configurations, or
who have not used the system in the past year to conduct a dry run of
their computer platform before the start of their observations.
- We continue monitoring the behavior of the new Vaisala relative
humidity and temperature measuring system.
- After the 2005 shutdown, TG installed a new guider similar to the
system in use at the 1.5m. He also implemented a new mode to allow use
of the guide camera over VNCviewer. WP and TG are implementing remote
- WP installed a new fiber optics run from the telescope to the
control room, which allows, e.g. for safe sensing of CCD temperatures.
- The telescope safety brakes remain on the list of backup projects.
- We are still exploring options for a replacement primary mirror.
- In July, Andy Szentgyorgyi, John Geary and Steve Amato completed
work in Cambridge on Keplercam, which replaces Minicam and the older
4shooter. It is a 4Kx4K single-chip 4-amplifier CCD with fast
readout, excellent sensitivity and cosmetics, and a long hold time in
its dewar. Its temperature is sensed and stored in image headers, thus
allowing observers to avoid problems with warm-ups. TG wrote the
control software, which was ready for the Keplercam first light in
late Jul 2005. It has performed extremely well since then.
- Stelircam continues to perform fairly well, although the red and
blue channels may not be available simultaneously. WP has continued to
address (so far) non-lethal failures in the electronics. A major
failure may imply the demise of the camera, because many of its
components are obsolete and no longer available.
Tillinghast 1.5-meter Telescope
The instruments currently available are FAST (dark time) and the Echelle
Telescope status and projects:
- We re-aluminized the primary and secondary mirrors during the Aug
2005 shutdown. We were able to use the "Sunnyside" facility at the U
of A again this year. The adhesion and reflectivity of the aluminum
- We monitor the sensitivity of FAST and the telescope optics by
observing a small set of standard stars with a 5" slit and the 300gpm
for details. Since the refurbishment of the FAST optics in Sep 2003,
the throughput decreased 20% (40%) at red (blue) wavelengths. We
attribute that to the degradation of the Al coating of the primary and
to contamination of the surface of the FAST corrector optics. The
latter were initially sol-gel coated and thus were impossible to clean
without losing the coating. After the Aug 2005 shutdown and mirror
re-aluminizations, we replaced the corrector with another with
hard-coated optics (see below for more on this corrector). These steps
along with periodic cleaning of the FAST optics have brought the
sensitivity back to the levels in Sep 2003, as desired.
- On Fri 3 June 2005, the dome control electronics failed during WP's
attempt to trace a dome problem. Because of another contemporaneous
failure, we had no spares, and as it was a weekend, replacement parts
arrived only the following Monday and the telescope was down for 3
- In Sep 2005, WP discovered that re-occurring dome problems were
mechanical, not electrical. The encoder-measuring wheel would not
contact the dome at certain points of rotation. The radial rollers
were at the extremes of their adjustment slots. WP loosened, adjusted
and tightened some of the rollers, and obtained the correct adjustment
procedure from Observadome. WP and RH adjusted the rest of the rollers
and the dome was then able to behave correctly.
- Beginning in Spring 2005, we had infrequent events where the
telescope would lock up in RA. There seemed to be no pattern related
to position or temperature. The gears were cleaned and re-greased as
the finer points of grease became clearer. The manufacturer of the
grease pointed out that we had used the wrong kind for many years; WP
promptly acquired the correct type. Springs were adjusted to apply
more or less pressure between the worm and the wormgear. The lock-ups
would disappear, but then reappear. Their frequency started increasing
in late 2005. During the night of 16 Jan 2006, the problem came to a
head, and the telescope was down for the following 2 nights. The
problem was finally pinpointed to the RA drive motor, which had
failed. We replaced it with a spare drive and sent out the failed
motor for refurbishment. After the replacement, significant
oscillations were apparent. During the following week, RH aligned and
repaired the flexible RA drive coupling and aligned the worm and the
wormgear, which eliminated the oscillations.
- RH and support worked on the dome, which had been cranky. An
insufficient number of the 32 axial rollers were making contact with
the base of the dome. After adding shims, about 70% of the rollers
make contact most of the time, which is near the design goal.
Instrument status and projects:
- At the end of the Aug 2004 shutdown, we replaced the sol-gel coated
corrector with a hard-coated version (optics were in hand and Warren
Brown assembled them). Unfortunately, this corrector produced
strangely flared stellar images on the guider. We reverted to the
sol-gel coated corrector and returned the other one to WB. The problem
was one of sligthly misaligned optics. Warren re-mounted the optics
and we swapped correctors again after shutdown in Aug 2005.
- We received the new FAST3 CCD with its own dewar and electronics in
early Jan 2006. TG had the requisite software ready at that time. On
Jan 12, we aligned and tested the new CCD on FAST. We discovered that
the noise level, e.g. on flats, was too high, and the sensitivity was
significantly below expectations. John Geary concurred, and requested
that we ship FAST3 back to Cambridge. With Steve Amato, he discovered
the cause was an incorrect "signal file" that was easily corrected.
They also repaired a faulty temperature sensor. We received FAST3 back
at FLWO on Feb 13. We plan to test it again on Mar 7, after the
current FAST run, and to bring it into service thereafter if its
performance is adequate.
- The Echelle continued producing useful data, in spite of its age
which causes occasional hiccups. As an example, one of these required
a repair to a controller that is no longer made. The manufacturer,
Oriel (now owned by Newport), claimed they could not repair it. WP
repaired it by building a new controller for it.
- The new high-throughput, fiber-fed TRES (Tillinghast Reflection
Echelle Spectrograph) construction is proceeding apace. Andy Szentgyorgyi is
working on this project with SAO pre-doc Gabor Furesz. We expect
delivery of TRES to FLWO in late Fall 2006.
PAIRITEL 1.3-meter Telescope
- The telescope is continuing nightly operations and is highly
- WP and RH designed and built an LN2 auto-refill system that
includes remote sensing of its status. The system has proven to be
- A small number of items remain for full automated operations:
remote control of the dome fan and power to the telescope drives.
- Contact Josh Bloom at Berkeley if you have a project that you would
like to have added to the queue.
The HATs (Hungarian Automated Telescopes) are a small autonomous
observatory designed for robotic observations
- Gaspar Bakos, now a Hubble Fellow at CfA, conducted regular
observations with his 3 FLWO HATs.
- GB added 2 new HATs at FLWO, one identical to the previous 3, and a
larger one, TopHAT (a 25-cm Baker Cassegrain), for a total of 5 HATs.
- The program to monitor sky brightness (see Reports No. 2-4) and
determine the effect of light pollution continues. We are currently
analyzing data from the past 2 years.
- The FLWO support crew added insulation to the HAT control room in
Apr 2005. It is essential to maintain acceptable temperatures now that
the number of HAT computers has increased significantly.
- Please refer to the MMT's own quarterly report.
- Mike Calkins and Perry Berlind assisted production observations as
fiber-Meisters and all-around observers with Hectospec and
Hectochelle. They also assisted with Megacam observations. That has
meant requesting that observers, especially graduate students, come to
FLWO to observe with FAST and the Echelle. We are thankful to the many
dedicated observers who have been willing to travel here. But we will
continue needing help; please consult the schedules and consider
coming to FLWO!
- Ridge staff continued providing assistance with SAO instrument and
computer installation and maintenance, as required. In particular, WH and
TG helped with network software and equipment, and RH built doors for
the new Hecto dark enclosure.
- On Jul 7 2005, lightning started a fire on Florida Peak, about 5
miles NE from FLWO. The fire grew to encompass about 23,000 acres when
it was 100% contained on July 24. On July 12 we evacuated the mountain
following a request from the Forest Service. FLWO was seriously
threatened, with the fire pushing to cross over Jack Mountain (between
Wrightson and Hopkins) toward the ridge on July 15-17. That put
flames within 0.5 miles or so of the ridge, but thanks to the efforts
of firefighters, especially the aerial crews that dropped retardant,
the fire was held at bay. A good combination of winds and, finally,
the start of the monsoon rains on Jul 17 started extinguishing the
fire. Additional downpours, typical of the monsoon, led to its full
demise by July 24. For additional details, visit
There was no damage to any equipment.
- We continue to update online documentation as changes occur, often
thanks to observers' comments. Please visit often!
- The issue of network safety remains a high priority. TG continues
monitoring the network. He also installed Spamassassin on all Ridge
- Dan Brocious and EF continued working on containing light
pollution. In Summer 2005, we participated in meetings in Bisbee,
where Cochise County was in the process of adopting a new outdoor
lighting code. The county adopted a good code, effective 02 Dec
2005. The final version included input from residents, amateur
astronomers and from us, including Grant Williams of the MMT and Liz
Alvarez from the U of A. We had a meeting in Jul 2005 with Santa
Cruz County staff who would enforce the code in the county. We expect
action on the code in 2006!
- In Apr/May 2005, two bills were narrowly approved by the AZ
legislature that could have allowed increases in light pollution, had
they not been vetoed by Governor Janet Napolitano. Senate Bill SB1193
would have weakened the ability of municipalities to enforce local
ordinances covering billboard and outdoor advertisements (including
their lighting). House Bill HB2461 would have allowed flashing LED
electronic billboards to be erected "off-premise" and was strongly
opposed by neighborhood associations and the astronomical community
and industry in Southern Arizona. FLWO joined the astronomical
community in AZ in this opposition.
Emilio E. Falco
670 Mt. Hopkins Road
P.O. Box 97
Amado, AZ 85645 USA